OpenStudy (anonymous):

3^2 + 6^2 = r^2 whats R?? im still stuggling

OpenStudy (31356):

radius

OpenStudy (31356):

Medal?

OpenStudy (31356):

If helped...

OpenStudy (destinymasha):

i think r is just the variable.....

OpenStudy (destinymasha):

used to represent the missing number lol

OpenStudy (31356):

It could be either

OpenStudy (destinymasha):

maybe

OpenStudy (mertsj):

3^2 = 9 6^2=36 9+36=45 so r^2=45 and r = sqrt45 or 3sq5

OpenStudy (anonymous):

i know it means radius, i mean like 3^2 (9) + 6^2 (36) = what number that i can eventualy write as 3^2+6^2=__^2

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

\[9+36=r^2\] \[45=r^2\] \[\sqrt{45}=\sqrt{r^2}\] \[3\sqrt{5}=r\]

OpenStudy (anonymous):

i know what i did wrong. its -3^2 which changes everything, right?

OpenStudy (mertsj):

Is it -3^2 or (-3)^2 ???

OpenStudy (anonymous):

(-3,6)

OpenStudy (mertsj):

I think the only sensible thing for you to do is post the ENTIRE problem.